What are two consecutive numbers with the sum of 95?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

What are two consecutive numbers with the sum of 95?
1st: x
2nd: x+1
---
Equation:
sum = 95
x+x+1 = 95
2x = 94
x = 47 (1st)
x+1 = 48 (2nd)
====================
Cheers,
Stan H.
====================

Answer by Alan3354(69443)

95/2 = 47.5
--> 47 & 48